# Privacy-Preserving Dataset Combination

### v1

@@ 4.2. Experimental Questions
4.3. Consistency Between Plaintext and Encrypted Computations
@@
For each source hospital $H_{o}$, use all full samples for $\mathcal{D}_{i}$. Between $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ and $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ on $\mathcal{D}_{o}$ and $\mathcal{D}_{i}$ for all remaining hospitals $H_{i}$, $\mathbb{E}_{H_{o}\sim\mathbf{H}}[\rho]=0.908$ with a range of $[0.691,1.0]$, obtaining $p<0.02$ across all hospitals. Between $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}}$ and $\mathrm{KL}_{\mathcal{X}}$, $\mathbb{E}_{H_{o}\sim\mathbf{H}}[\rho]=0.9303$ with a range of $[0.455,0.991]$, with 11 of 12 hospitals achieving p-values below $0.05$. After applying Hochberg false discovery rate correction (Benjamini1995ControllingTF, ), our p-values remain significant. This range is an artifact of sweeping hyperparameters independently in plaintext and encrypted optimisations. When we unify the SGD hyperparameters, we indeed get a tighter range. For all 12 hospitals, see appended Appendix C for details.
@@
We further simulate the effect by adding encryption through its impact on the downstream AUC. This examines whether there will be a shift in the full hospital ranking, if we switch from a plaintext setup to encrypted. For $H_{o}\sim\mathbf{H}$, we measure $\delta_{i}$ that results from adding $\mathcal{D}_{i}$ to $\mathcal{D}_{o}$ for all $i$. This correlates all target hospitals $\{H_{i}\}$ with their ground truths $\{\delta_{i}\}$ in the case of picking a single target hospital. We find the linear coefficient for encrypted $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ to be $-0.182$ and plaintext $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ to be $-0.184$ ($99\%$ matching). Both $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}}$ and $\mathrm{KL}_{\mathcal{X}}$ have a linear coefficient of $-0.164$ with $\delta_{i}$. For all strategies’ correlations with ground truth at $n=1$, see Appendix A.
Figure 2. AUC change $\delta_{T}$, after including top-$3$ hospital per strategy, over all source hospitals. Our private strategy $\pi_{p}$ is compared with (a) demographic-based $\pi_{d}$ for gender, age, and race. (b) plain text limited-sample $\pi_{s}(k=K)$ for $k=300$ (10%) and $k=30$ (1%).
4.4. Positivity of SecureKL and Practical Implications
@@
4.4.1. Comparing With Alternative Strategies
@@ 4.4.2. SecureKL Analysis
In the case where that output can be sensitive, i.e., when hospitals query each other multiple times and accrue information through the score function, the output can also be made privacy-preserving through differentially private data releases, such as using randomized response (dwork2014algorithmic, ).
In theory, any data combination method (if Turing-complete) can be made private; yet, in practice, balancing the right trade-off of utility and privacy is non-trivial. Barring engineering difficulties, not all algorithms readily adapt efficiently in private. Prior work  (xu2022data, ) included the trained model and test data in private; while relatively exact, complex methods would exacerbate the same operational limitations discussed in Section 4.5.3.
@@
4.5. Error Analysis
4.5.1. Underlying Score Limitations
@@
4.5.2. Sometimes, Not All Underlying Data Is Needed
@@
Figure 3. Hospital 243: Underlying KL-score performs poorly
[…]
NEED UPDATING some conclusion and description
Figure 4. Hospital 420: Underlying KL-score performs well
[…]
NEED UPDATING
Figure 5. Left: $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ outperforms $\pi_{p}(k=30)$ and $\pi_{0}$. Middle: All strategies perform similarly. Right: $\pi_{p}(k=300)$ outperforms $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$. In all panels, the bars represent the standard deviation.
[…]
NEED UPDATING
4.5.3. Encrypted Computation Limitations
@@
  4. (4)
    Inherent to Secure Computation: When the method requires significant hyper-parameter tuning, such as using SGD on small batch data with learning rate schedules, plaintext tuning may not transfer perfectly. As detailed in Appendix  B, our hyperparameters for SGD are indeed different in encrypted and plaintext settings. However, encrypted computation hides loss curves and training details by default, complicating development.
5. Related Work
5.1. Data Valuation and Pricing
The question of how to assess the impact and therefore worth of data has been well studied. Data valuation as a field seeks to quantify the contribution of individual data points or datasets to model performance. Shapley value-based approaches provide theoretically grounded valuations but scale poorly to large datasets (ghorbani2019data, ). More efficient methods include influence functions (koh2017understanding, ) and leave-one-out testing (cook1979influential, ). Recent work has extended these concepts to dataset-level valuation (jia2019towards, ). As practitioners create larger datasets, the emergence of data marketplaces has sparked interest in data pricing mechanisms (kumar2020marketplace, ). Query-based pricing (koutris2015query, ) and outcome-driven valuations (radic2024pricing, ) aim to balance seller compensation with buyer utility. While these approaches inform fair data exchange, they typically assume direct access to data, unlike our privacy-preserving method.
5.2. Alternative Approaches to Data Sharing
@@
Synthetic data generation has emerged as a promising approach to expand training datasets while preserving privacy. Generative adversarial networks (GANs) have shown success in generating realistic cancer incidence data (goncalves2020generation, ), medical imaging data (thambawita2022singan, ), and electronic health records (baowaly2019synthesizing, ). These methods can preserve statistical properties of the original data while providing differential privacy guarantees. Transforming data into a similar form that de-sensitizes certain attributes can be desirable (drechsler2011synthetic, ; howe2017synthetic, ; nikolenko2021synthetic, ; gonzales2023synthetic, ; sweeney2002k, ). Yet, to still preserve the utility of the dataset transformed for analytics or learning tasks is challenging by itself (jordon2021hide, ). Additionally, outside of the scope of sensitive data that is transformed, little privacy guarantee is available, leading to re-identification risks (narayanan2006break, ; jordon2021hide, ).
However, evaluation of synthetic medical data reveals challenges in capturing rare conditions and maintaining consistent relationships between multiple health variables (goncalves2020generation, ). For tabular data, methods like CTGAN (xu2019modeling, ) and TVAE (xu2019modeling, ) have demonstrated ability to learn complex distributions while preserving correlations between features. However, these approaches often struggle with high-dimensional data and can introduce subtle biases that impact downstream model performance (assefa2020generating, ). Recent work has also explored combining synthetic data with differential privacy to provide formal privacy guarantees (jordon2018pate, ). While these methods offer stronger privacy protection, they often face significant utility loss, particularly for rare but important cases in the original dataset (yang2024tabular, ).
@@
Transfer learning via public pretraining has become increasingly popular for domains with limited private data access. BioBERT (lee2020biobert, ) and ClinicalBERT (alsentzer2019publicly, ) demonstrated that pretraining on PubMed abstracts and clinical notes can improve performance on downstream medical tasks. Similar approaches have emerged in other regulated domains, including FinBERT (liu2021finbert, ) for financial applications. However, the effectiveness of transfer learning depends heavily on domain alignment. One study showed that continued pretraining on domain-specific data significantly outperforms generic pretraining when domains differ substantially (gururangan2020don, ). This presents challenges for highly specialized fields where public data may not capture domain-specific patterns (le2022language, ). Recent work has explored methods to quantify and optimize domain adaptation. Adaptive pretraining strategies (howard2018universal, ) and domain-specific vocabulary augmentation (gururangan2020don, ) have shown promise in bridging domain gaps. However, these approaches still require substantial compute and may not fully capture specialized domain knowledge present in private datasets.
5.3. Benefits of Data Scaling Beyond Performance
Recent work has demonstrated that increasing dataset size and diversity yields benefits beyond raw performance metrics. Large-scale training data has been shown to improve model robustness to distribution shifts (taori2020measuring, ) and reduce demographic performance disparities (chen2018my, ). Studies of vision models trained on increasingly large datasets show improved out-of-distribution generalization (miller2021accuracy, ). Similarly, language models trained on diverse data demonstrate better performance across different domains and demographic groups (dodge2021documenting, ).
5.4. Secure and Confidential Computation
@@
FHE enables arbitrary additions and multiplications on encrypted inputs. While it represents the gold standard for encrypted computation, adapting it to modern machine learning is challenging due to computational constraints from growing ciphertext size. FHE implementations typically use lattice-based schemes requiring periodic “bootstrapping” (key refreshing and noise reduction through re-encryption) via methods like the CKKS scheme (cheon2017homomorphic, ). This introduces cryptographic parameters that non-experts struggle to configure effectively.
@@
SMPC enables multiple parties to compute functions over private inputs while revealing only the final output (yao1982protocols, ; shamir1979share, ). An MPC system uses key exchanges, encryption schemes, and secure communication to ensure only encrypted data leaves owner control (micali1987play, ). Though generally faster than FHE, SMPC’s engineering complexity and communication overhead can limit adoption. Traditional private training approaches that completely hide data can also impede essential model development tasks like inspection, monitoring, and debugging.
5.5. Secure Data Combination
Recent work has explored methods for securely combining datasets while preserving privacy and improving model performance. Early approaches focused on using secure multi-party computation to enable multiple parties to jointly train models without sharing raw data (aono2017privacy, ). However, these methods often struggled with computational overhead and communication costs when dealing with large-scale datasets (mcmahan2017communication, ). More recent techniques have introduced frameworks for evaluating potential data partnerships before commitment. These approaches use privacy-preserving protocols to estimate the compatibility and complementarity of different datasets (leung2019towards, ; chakraborty2024privacy, ). Some methods focus specifically on measuring distribution shifts between datasets without revealing sensitive information (duan2021flexible, ). Several systems have been developed to facilitate secure data combination in specific domains. In healthcare, methods have been proposed for securely combining patient records across institutions while maintaining HIPAA compliance (raisaro2018m, ). Financial institutions have explored similar approaches for combining transaction data while preserving client confidentiality (liu2021finbert, ).
5.6. Other Privacy-Preserving Methods
@@
Cross-silo federated, decentralized, and collaborative machine learning  (mcmahan2017communication, ; li2020federated, ; bonawitztowards, ; kairouz2021advances, ) focus on acquiring more data through improved data governance and efficient system design. Healthcare machine learning is considered especially suitable, as health records are often isolated (rieke2020future, ; xu2021federated, ; nguyen2022federated, ). Yet, even though no raw data is shared, model parameters or gradients flow through the system. As the federated computing paradigm offer no privacy guarantee, the system is vulnerable to model inversion (geiping2020inverting, ) and gradients leakage attacks (boenisch2023curious, ; zhu2019deep, ). A subtle but urgent concern is that privacy risks discourage the very formation of the federation when optimisation is traded off with privacy (lyu2022privacy, ; raynal2024conflict, ). Building on the insight that useful data is often disparately owned, we tackle the specific incentive problem between pairs of data players where one side trains the model, instead of scaling up a federation (number of parties) to address data access issues. We thus focusing on making this exchange efficient, accurate, and private.
Compared to vanilla Federated Learning, an MPC system (shamir1979share, ; yao1982protocols, ; bonawitz2017practical, ; knott2021crypten, ) provides stronger guarantee in terms of input security. Model owners and data owners can potentially federate their proprietary data, including model weights, training, and testing data, can work together under stringent privacy requirements. Our work extends the line of works  (xu2022data, ; yang2024fedfed, ; bonawitz2017practical, ) that demonstrates the potential of incorporating MPC in various federated scenarios. On the practical side, unlike mobile-based networks for secure federated learning protocols (bonawitztowards, ), our system assumes a smaller number of participants, where communication cost and runtime are not dominant concerns.
@@
Differential privacy (DP) (dwork2006differential, ) offers formal privacy guarantees for sharing data and training machine learning models. While DP mechanisms can protect individual privacy when releasing model outputs or aggregated statistics, they face significant limitations for inter-organizational data sharing. The primary challenge is that DP operates on already-pooled data, but organizations are often unwilling to share their raw data in the first place (dwork2014algorithmic, ). Even when organizations are willing to share data, the privacy guarantees of DP come at a substantial cost to utility, particularly in machine learning applications. DP-SGD, the standard approach for training deep neural networks with differential privacy, significantly degrades model performance compared to non-private training (abadi2016deep, ). This performance impact is especially pronounced in data-constrained settings, where recent work has shown that large models rely heavily on memorization of rare examples that DP mechanisms tend to obscure (feldman2020neural, ). The privacy-utility trade-off becomes even more challenging when dealing with high-dimensional data or complex learning tasks. Studies have demonstrated that achieving meaningful privacy guarantees while maintaining acceptable model performance requires prohibitively large datasets (bagdasaryan2020backdoor, ). This limitation is particularly problematic in specialized domains like healthcare, where data is inherently limited and performance requirements are stringent (geyer2017differentially, ). Recent work has attempted to improve the privacy-utility trade-off through advanced composition theorems and adaptive privacy budget allocation (papernot2021tempered, ). However, these approaches still struggle to match the performance of non-private training, especially when working with modern deep learning architectures (tramer2020differentially, ). While differential privacy offers important theoretical guarantees, our work focuses on the practical challenge of enabling data owners to evaluate potential partnerships before sharing any data, addressing a key barrier to collaboration that DP alone cannot solve.
6. Conclusion

### v2

@@ 4.2 Experimental Questions
4.3 Consistency Between Plaintext and Encrypted Computations
@@
For each source hospital $H_{o}$, use all full samples for $\mathcal{D}_{i}$. Between $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ and $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ on $\mathcal{D}_{o}$ and $\mathcal{D}_{i}$ for all remaining hospitals $H_{i}$, $\mathbb{E}_{H_{o}\sim\mathbf{H}}[\rho]=0.908$ with a range of $[0.691,1.0]$, obtaining $p<0.02$ across all hospitals. Between $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}}$ and $\mathrm{KL}_{\mathcal{X}}$, $\mathbb{E}_{H_{o}\sim\mathbf{H}}[\rho]=0.9303$ with a range of $[0.455,0.991]$, with 11 of 12 hospitals achieving p-values below $0.05$. After applying Hochberg false discovery rate correction [42], our p-values remain significant. This range is an artifact of sweeping hyperparameters independently in plaintext and encrypted optimisations. When we unify the SGD hyperparameters, we indeed get a tighter range. For all 12 hospitals, see appended Appendix 9 for details.
@@
We further simulate the effect by adding encryption through its impact on the downstream AUC. This examines whether there will be a shift in the full hospital ranking, if we switch from a plaintext setup to encrypted. For $H_{o}\sim\mathbf{H}$, we measure $\delta_{i}$ that results from adding $\mathcal{D}_{i}$ to $\mathcal{D}_{o}$ for all $i$. This correlates all target hospitals $\{H_{i}\}$ with their ground truths $\{\delta_{i}\}$ in the case of picking a single target hospital. We find the linear coefficient for encrypted $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ to be $-0.182$ and plaintext $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ to be $-0.184$ ($99\%$ matching). Both $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}}$ and $\mathrm{KL}_{\mathcal{X}}$ have a linear coefficient of $-0.164$ with $\delta_{i}$. For all strategies’ correlations with ground truth at $n=1$, see Appendix 7.
Figure 2: AUC change $\delta_{T}$, after including top-$3$ hospital per strategy, over all source hospitals. Our private strategy $\pi_{p}$ is compared with (a) demographic-based $\pi_{d}$ for gender, age, and race. (b) plain text limited-sample $\pi_{s}(k=K)$ for $k=300$ (10%) and $k=30$ (1%).
4.4 Positivity of SecureKL and Practical Implications
@@
4.4.1 Comparing With Alternative Strategies
@@ 4.4.2 SecureKL Analysis
In the case where that output can be sensitive, i.e., when hospitals query each other multiple times and accrue information through the score function, the output can also be made privacy-preserving through differentially private data releases, such as using randomized response [43].
In theory, any data combination method (if Turing-complete) can be made private; yet, in practice, balancing the right trade-off of utility and privacy is non-trivial. Barring engineering difficulties, not all algorithms readily adapt efficiently in private. Prior work  [44] included the trained model and test data in private; while relatively exact, complex methods would exacerbate the same operational limitations discussed in Section 4.5.3.
@@
4.5 Error Analysis
4.5.1 Underlying Score Limitations
@@
4.5.2 Sometimes, Not All Underlying Data Is Needed
@@
Figure 3: Hospital 243: Underlying KL-score performs poorly
Figure 4: Hospital 420: Underlying KL-score performs well
Figure 5: Left: $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ outperforms $\pi_{p}(k=30)$ and $\pi_{0}$. Middle: All strategies perform similarly. Right: $\pi_{p}(k=300)$ outperforms $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$. In all panels, the bars represent the standard deviation.
4.5.3 Encrypted Computation Limitations
@@
  4. 4.
    Inherent to Secure Computation: When the method requires significant hyper-parameter tuning, such as using SGD on small batch data with learning rate schedules, plaintext tuning may not transfer perfectly. As detailed in Appendix  8, our hyperparameters for SGD are indeed different in encrypted and plaintext settings. However, encrypted computation hides loss curves and training details by default, complicating development.
5 Related Work
5.1 Data Valuation and Pricing
The question of how to assess the impact and therefore worth of data has been well studied. Data valuation as a field seeks to quantify the contribution of individual data points or datasets to model performance. Shapley value-based approaches provide theoretically grounded valuations but scale poorly to large datasets [45]. More efficient methods include influence functions [46] and leave-one-out testing [47]. Recent work has extended these concepts to dataset-level valuation [48]. As practitioners create larger datasets, the emergence of data marketplaces has sparked interest in data pricing mechanisms [49]. Query-based pricing [50] and outcome-driven valuations [51] aim to balance seller compensation with buyer utility. While these approaches inform fair data exchange, they typically assume direct access to data, unlike our privacy-preserving method.
5.2 Alternative Approaches to Data Sharing
@@
Synthetic data generation has emerged as a promising approach to expand training datasets while preserving privacy. Generative adversarial networks (GANs) have shown success in generating realistic cancer incidence data [52], medical imaging data [53], and electronic health records [54]. These methods can preserve statistical properties of the original data while providing differential privacy guarantees. Transforming data into a similar form that de-sensitizes certain attributes can be desirable [55, 56, 57, 58, 59]. Yet, to still preserve the utility of the dataset transformed for analytics or learning tasks is challenging by itself [60]. Additionally, outside of the scope of sensitive data that is transformed, little privacy guarantee is available, leading to re-identification risks [61, 60].
However, evaluation of synthetic medical data reveals challenges in capturing rare conditions and maintaining consistent relationships between multiple health variables [52]. For tabular data, methods like CTGAN [62] and TVAE [62] have demonstrated ability to learn complex distributions while preserving correlations between features. However, these approaches often struggle with high-dimensional data and can introduce subtle biases that impact downstream model performance [63]. Recent work has also explored combining synthetic data with differential privacy to provide formal privacy guarantees [64]. While these methods offer stronger privacy protection, they often face significant utility loss, particularly for rare but important cases in the original dataset [65].
@@
Transfer learning via public pretraining has become increasingly popular for domains with limited private data access. BioBERT [17] and ClinicalBERT [15] demonstrated that pretraining on PubMed abstracts and clinical notes can improve performance on downstream medical tasks. Similar approaches have emerged in other regulated domains, including FinBERT [66] for financial applications. However, the effectiveness of transfer learning depends heavily on domain alignment. One study showed that continued pretraining on domain-specific data significantly outperforms generic pretraining when domains differ substantially [16]. This presents challenges for highly specialized fields where public data may not capture domain-specific patterns [67]. Recent work has explored methods to quantify and optimize domain adaptation. Adaptive pretraining strategies [68] and domain-specific vocabulary augmentation [16] have shown promise in bridging domain gaps. However, these approaches still require substantial compute and may not fully capture specialized domain knowledge present in private datasets.
5.3 Benefits of Data Scaling Beyond Performance
Recent work has demonstrated that increasing dataset size and diversity yields benefits beyond raw performance metrics. Large-scale training data has been shown to improve model robustness to distribution shifts [23] and reduce demographic performance disparities [69]. Studies of vision models trained on increasingly large datasets show improved out-of-distribution generalization [14]. Similarly, language models trained on diverse data demonstrate better performance across different domains and demographic groups [70].
5.4 Secure and Confidential Computation
@@
FHE enables arbitrary additions and multiplications on encrypted inputs. While it represents the gold standard for encrypted computation, adapting it to modern machine learning is challenging due to computational constraints from growing ciphertext size. FHE implementations typically use lattice-based schemes requiring periodic “bootstrapping” (key refreshing and noise reduction through re-encryption) via methods like the CKKS scheme [71]. This introduces cryptographic parameters that non-experts struggle to configure effectively.
@@
SMPC enables multiple parties to compute functions over private inputs while revealing only the final output [33, 34]. An MPC system uses key exchanges, encryption schemes, and secure communication to ensure only encrypted data leaves owner control [72]. Though generally faster than FHE, SMPC’s engineering complexity and communication overhead can limit adoption. Traditional private training approaches that completely hide data can also impede essential model development tasks like inspection, monitoring, and debugging.
5.5 Secure Data Combination
Recent work has explored methods for securely combining datasets while preserving privacy and improving model performance. Early approaches focused on using secure multi-party computation to enable multiple parties to jointly train models without sharing raw data [73]. However, these methods often struggled with computational overhead and communication costs when dealing with large-scale datasets [74]. More recent techniques have introduced frameworks for evaluating potential data partnerships before commitment. These approaches use privacy-preserving protocols to estimate the compatibility and complementarity of different datasets [75, 76]. Some methods focus specifically on measuring distribution shifts between datasets without revealing sensitive information [77]. Several systems have been developed to facilitate secure data combination in specific domains. In healthcare, methods have been proposed for securely combining patient records across institutions while maintaining HIPAA compliance [78]. Financial institutions have explored similar approaches for combining transaction data while preserving client confidentiality [66].
5.6 Other Privacy-Preserving Methods
@@
Cross-silo federated, decentralized, and collaborative machine learning  [74, 79, 80, 81] focus on acquiring more data through improved data governance and efficient system design. Healthcare machine learning is considered especially suitable, as health records are often isolated [82, 83, 84]. Yet, even though no raw data is shared, model parameters or gradients flow through the system. As the federated computing paradigm offer no privacy guarantee, the system is vulnerable to model inversion [85] and gradients leakage attacks [86, 87]. A subtle but urgent concern is that privacy risks discourage the very formation of the federation when optimisation is traded off with privacy [88, 89]. Building on the insight that useful data is often disparately owned, we tackle the specific incentive problem between pairs of data players where one side trains the model, instead of scaling up a federation (number of parties) to address data access issues. We thus focusing on making this exchange efficient, accurate, and private.
Compared to vanilla Federated Learning, an MPC system [34, 33, 90, 35] provides stronger guarantee in terms of input security. Model owners and data owners can potentially federate their proprietary data, including model weights, training, and testing data, can work together under stringent privacy requirements. Our work extends the line of works  [44, 91, 90] that demonstrates the potential of incorporating MPC in various federated scenarios. On the practical side, unlike mobile-based networks for secure federated learning protocols [80], our system assumes a smaller number of participants, where communication cost and runtime are not dominant concerns.
@@
Differential privacy (DP) [92] offers formal privacy guarantees for sharing data and training machine learning models. While DP mechanisms can protect individual privacy when releasing model outputs or aggregated statistics, they face significant limitations for inter-organizational data sharing. The primary challenge is that DP operates on already-pooled data, but organizations are often unwilling to share their raw data in the first place [43]. Even when organizations are willing to share data, the privacy guarantees of DP come at a substantial cost to utility, particularly in machine learning applications. DP-SGD, the standard approach for training deep neural networks with differential privacy, significantly degrades model performance compared to non-private training [93]. This performance impact is especially pronounced in data-constrained settings, where recent work has shown that large models rely heavily on memorization of rare examples that DP mechanisms tend to obscure [94]. The privacy-utility trade-off becomes even more challenging when dealing with high-dimensional data or complex learning tasks. Studies have demonstrated that achieving meaningful privacy guarantees while maintaining acceptable model performance requires prohibitively large datasets [95]. This limitation is particularly problematic in specialized domains like healthcare, where data is inherently limited and performance requirements are stringent [96]. Recent work has attempted to improve the privacy-utility trade-off through advanced composition theorems and adaptive privacy budget allocation [97]. However, these approaches still struggle to match the performance of non-private training, especially when working with modern deep learning architectures [98]. While differential privacy offers important theoretical guarantees, our work focuses on the practical challenge of enabling data owners to evaluate potential partnerships before sharing any data, addressing a key barrier to collaboration that DP alone cannot solve.
6 Conclusion